### v1

# Infinity-Parser: Layout-Aware Reinforcement Learning for Scanned Document Parsing

Abstract
Automated parsing of scanned documents into richly structured, machine-readable formats remains a critical bottleneck in Document AI, as traditional multi-stage pipelines suffer from error propagation and limited adaptability to diverse layouts. We introduce layoutRL, an end-to-end reinforcement-learning framework that trains models to be explicitly layout-aware by optimizing a composite reward of normalized edit-distance, paragraph-count accuracy, and reading-order preservation. Leveraging our newly released dataset—Infinity-Doc-55K, which combines 55K high-fidelity synthetic scanned document parsing data with expert-filtered real-world documents—we instantiate layoutRL in a vision-language-model–based parser called Infinity-Parser. Evaluated on English and Chinese benchmarks for OCR, table and formula extraction, and reading-order detection, Infinity-Parser achieves new state-of-the-art performance in both accuracy and structural fidelity, outpacing specialist pipelines and general-purpose vision-language models. We will publicly release our code and dataset to accelerate progress in robust document understanding.
1 Introduction
Automating the parsing of scanned documents into layout, machine-readable formats is a longstanding challenge in Document AI [19, 52, 59, 62, 69]. While scanned documents provide fine-grained control over visual layouts, they lack explicit semantic markers for essential elements such as paragraphs, headings, tables, and lists. This absence introduces structural ambiguity, making it difficult to recover the hierarchical organization of documents and limiting the scalability of traditional rule-based or multi-stage parsing systems [52, 37]. Conventional approaches typically rely on a sequence of supervised tasks—including layout detection, optical character recognition (OCR), table recognition and math formula recognition —followed by heuristic or scripted post-processing to reconstruct document layout [6, 61, 25, 59, 2, 8]. These pipelines not only demand extensive annotated resources but also suffer from error propagation and limited adaptability to diverse or unseen layouts.
Recent advances in vision-language models (VLMs) have enabled end-to-end approaches that directly map visual document inputs to layout outputs such as Markdown [3, 8, 63, 48]. However, off-the-shelf pretraining objectives—focused on image–text alignment or language modeling—do not inherently encourage preserving a document’s spatial arrangement or reading order. These approaches simplify traditional multi-stage pipelines and offer new possibilities for document parsing, existing models still struggle to generalize effectively on downstream tasks [52, 59]. A key reason lies in the fundamental mismatch between pretraining objectives and the structural understanding required for document parsing—most pretraining focuses on image-text alignment and reading comprehension, without optimizing for the structural complexities inherent in documents. Compared to standard text, documents exhibit richer, more intricate layouts, characterized by hierarchical layouts, nested elements, and multimodal content [17, 44, 65, 64]. This complexity introduces challenges that conventional token-level supervision cannot address: prediction errors on a few critical tokens can propagate into substantial structural deviations, and current methods lack fine-grained, layout-aware signals to guide models toward outputs that are both semantically accurate and structurally faithful [11, 24, 42]. Recent studies show that reinforcement learning (RL) achieves stronger generalization in rule-based and visual reasoning tasks [9]. Existing RL approaches remain constrained by coarse-grained, binary outcome rewards that fail to provide the fine-grained, layout-aware supervision needed for modeling complex document layouts [29, 56, 73, 32]. Designing effective layout-aware reward functions tailored for document parsing remains an underexplored challenge.
Figure 1: Overview of the layout-aware reinforcement learning framework. It includes data preparation, reinforcement finetuning, and reward calculation based on edit distance, count, and order to enhance structural understanding.
To address these limitations, we propose layoutRL, the first end-to-end reinforcement-learning framework that makes document parsers explicitly layout-aware, and construct a new dataset, as shown in Figure 1. Specially, first, to overcome the generalization gap caused by the mismatch between pretraining objectives and structural understanding, we incorporate verifiable rewards [43, 24, 42, 66] that provide document-level supervision beyond token-level signals, encouraging the model to learn transferable structural representations. Second, to mitigate structural hallucination under complex layouts, we design Edit Distance Reward [21] and Layout Parsing Reward signals, including count and order that enforce fine-grained alignment between predictions and ground-truth layouts, improving layout coherence. Additionally, to train our RL model we construct Infinity-Doc-55K, a 55 K-document corpus providing large-scale, high-quality supervision. It combines (1) high-fidelity synthetic scanned document parsing data—generated via HTML templates and browser rendering—and (2) expert-filtered real-world samples, pseudo-labeled through a cross-model agreement pipeline to capture genuine layout diversity. Finally, we train a vision-language-model–based parser, Infinity-Parser, achieves state-of-the-art results across English and Chinese benchmarks—covering OCR, table and formula extraction, and reading-order detection—and consistently outperforms both specialist pipelines and leading vision-language models.
We make the following contributions:
  • We propose layoutRL, a new reinforcement learning framework for end-to-end scanned document parsing, which explicitly trains models to be layout-aware by optimizing verifiable, multi-aspect rewards. Our multi-aspect reward design combines normalized edit distance, paragraph count accuracy, and reading order preservation, enabling stable training and improved structural robustness.
  • We introduce Infinity-Doc-55K, a large-scale dataset of 55,066 scanned documents that combines high-quality synthetic data with diverse real-world samples. The dataset features rich layout variations and comprehensive structural annotations, enabling robust training of document parsing models.
  • We train a VLM based model, Infinity-Parser, which sets new state-of-the-art performance across English and Chinese benchmarks for OCR, table and formula extraction, and reading-order detection—demonstrating substantial gains in both structural fidelity and semantic accuracy over specialist pipelines and general-purpose vision-language models.
Benchmark Document Domain Annotation Type End-to-End Task Exactly Match
BBox Text Table Formula Attributes OCR TR MFR ROD
End-to-end Eval Benchmarks
Fox [25] 2
Nougat [6] 1
GOT OCR 2.0 [59] 2
OmniDocBench [36] 9
End-to-end Train Dataset
DocStruct4M [13] -
olmOCR-mix [39] -
Infinity-Doc-55K 7
Table 1: A comparison between Infinity-Doc-55K and existing datasets. BBox: Bounding boxes. Text: Text in Unicode. Table: Table in LaTeX/HTML/Markdown. Formula: Formula in LaTeX. Attributes: Page- and BBox-Level Attributes. OCR: Optical Character Recognition; TR: Table Recognition; MFR: Math Formula Recognition; ROD: Reading Order Detection. Multi-Type Doc: Whether the dataset includes documents from multiple domains or categories.
2 Methodology
In this section, we first introduce Infinity-Doc-55K, our large-scale multimodal dataset for end-to-end scanned document parsing. We then describe our rule-based multi-aspect reward framework, which integrates edit distance, paragraph count, and order criteria under a unified reinforcement learning objective optimized via Group Relative Policy Optimization (GRPO).
2.1 Infinity-Doc-55K and Generation Pipelines
We introduce Infinity-Doc-55K, a large-scale, multimodal dataset of 55,066 richly annotated documents for end-to-end scanned document parsing. Unlike prior benchmarks that target isolated subtasks (e.g., layout detection, OCR, or table recognition), Infinity-Doc-55K provides holistic supervision by pairing rendered scanned documents pages with their ground-truth Markdown representations. This design enables training and evaluating models that directly translate visual inputs to layout outputs without relying on brittle, multi-stage pipelines. As shown in Table 1, compared to existing works, Infinity-Doc-55K not only significantly enhances task diversity but also substantially improves overall data quality through our proposed synthetic generation mechanism. More details on the data distribution are provided in the Data Details section of the Appendix.
Figure 2: Data construction pipelines for document parsing. (a) Real-world pipeline enhance quality by combining multiple expert models and layout analysis, yielding better-aligned supervision through intersection and reading order reasoning. (b) Synthetic pipeline leverages structured HTML templates and browser rendering to generate clean, exactly-aligned scanned document parsing data, ensuring high-quality supervision for end-to-end parsing.
To construct Infinity-Doc-55K, we design a dual-pipeline framework that integrates both synthetic and real-world document generation, as illustrated in Figure 2. This design addresses a critical limitation of traditional data construction pipelines, which often rely on weak supervision and pseudo-labeling from a single model applied to crawled scanned documents. These pipelines frequently suffer from noisy, misaligned, or incomplete annotations, especially in complex layouts or multilingual content, thus hindering model performance and generalization. To overcome these issues, our dual-pipeline framework is motivated by the need to balance annotation quality and structural diversity. The synthetic branch provides highly accurate, clean, and precisely aligned annotations at scale, while the real-world branch introduces naturally occurring layout variability and semantic richness, which are essential for building models that generalize robustly in practical applications.
Real-World Data We develop a real-world data construction pipeline to capture the structural complexity and natural layout variability of documents across practical domains. We collect diverse scanned documents from sources such as financial reports, medical records, academic papers, books, magazines, and web pages, covering both dense and sparse content layouts. To generate structural annotations, we adopt a multi-expert strategy, where specialized models handle different structural elements, such as layout blocks, texts, formulas, and tables. For example, overall layouts are analyzed by visual layout model [18], formula regions are processed by dedicated formula recognition model [50], tables are parsed by transformer-based table extractor [6]. A cross-validation mechanism is then applied to filter out inconsistencies by comparing the outputs of expert models and VLM [1]. Only regions with consistent predictions across models are retained as high-confidence pseudo-ground-truth annotations. This layout-aware filtering results in a rich and reliable dataset that reflects the complexity of real-world documents and supports robust document parsing model training.
Synthetic Data We design a synthetic data construction pipeline. We collect text and images from sources such as Wikipedia, web crawlers, and online corpora, and use Jinja [34] templates to inject sampled content into predefined single-, double-, or triple-column HTML layouts. These pages are rendered into scanned documents using a browser engine, followed by automated filtering to remove low-quality or overlapping images. Ground-truth annotations are extracted by parsing the original HTML to produce aligned Markdown representations. This synthetic approach not only significantly reduces construction costs and ensures annotation accuracy and structural diversity, but more importantly, it addresses the longstanding issue of imprecise or inconsistent supervision commonly found in pseudo-labeled datasets, providing high-quality and well-aligned supervision for training end-to-end models.
Figure 3: This figure illustrates a diverse collection of PDF document types commonly encountered in Infinity-Doc-55K, grouped into seven categories: Medical Reports, Synthetic Documents, Financial Reports, Academic Papers, Books, Magazines, and Web Pages.
Data Details As illustrated in Figure 3, the dataset spans seven diverse document domains, making it one of the most richly annotated and structurally varied resources to date. Each domain is represented by two sample pages, highlighting the broad variability in layout design, content structure, and semantic density. For instance, Medical Reports typically contain structured tables with clinical measurements and diagnostic notes. Synthetic Documents are algorithmically generated to replicate real-world formats, providing layout diversity for training robust parsers. Financial Reports feature dense tables and formal accounting records, while Academic Papers often follow two-column layouts with references, equations, and figures. Books combine narrative content with visual illustrations, and Magazines blend images and stylized text for reader engagement. Finally, Web Pages, when saved as PDFs, preserve HTML-based structures that integrate tables, lists, and dynamic elements. This visual taxonomy exemplifies the structural and semantic diversity present in real-world documents, highlighting the core challenge faced by document AI systems: reliably parsing heterogeneous layouts and extracting structured information across a wide variety of formats.
Figure 4: Overview of the Infinity-Parser training framework. Our model is optimized via reinforcement fine-tuning with edit distance, layout, and order-based rewards. The detailed formulations of these reward functions are provided in Section 2.2.
Table 13 provides an overview of the document types included in the Infinity-Doc-55K dataset, detailing the composition across both real-world and synthetic sources. The real-world portion consists of 48.6k documents spanning six domains: financial reports, medical reports, academic papers, books, magazines, and web pages. These documents were collected from the web and annotated using a pseudo-labeling pipeline based on expert model agreement. While this approach enables large-scale data acquisition, the resulting label quality is relatively low due to occasional inconsistencies across models. Additionally, real-world data collection incurs a high cost, especially in terms of manual filtering, formatting normalization, and layout validation.
2.2 RL with Layout-Aware Rewards
As illustrated in Figure 4, we employ a RL framework to directly optimize scanned document parsers, aiming to enhance both structural fidelity and semantic accuracy. To achieve this, we utilize GRPO [42], which enables learning from rule-based reward signals without relying on absolute values.
GRPO operates by generating a set of candidate Markdown outputs for each document and evaluating them using a multi-aspect reward, denoted as $R_{\text{Multi-Aspect}}$, which integrates multiple rule-based criteria into a unified supervisory signal. These raw rewards are then converted into relative advantage scores by comparing each candidate against others within the same group. This relative evaluation promotes training stability and encourages the selection of higher-quality outputs, eliminating the need for a learned value function or critic.
The multi-aspect reward $R_{\text{Multi-Aspect}}$ consists of three complementary components, each capturing a different aspect of parsing quality:
1. Edit Distance Reward ($R_{\mathrm{dist}}$)
We define the edit distance reward based on the normalized Levenshtein distance $D(y,\hat{y})$ between the predicted output $\hat{y}$ and the reference output $y$:
$R_{\text{dist}}=1-\frac{D(y,\hat{y})}{\max(N,M)},$
where $N=|y|$ and $M=|\hat{y}|$ are the lengths of the reference and predicted sequences, respectively. The distance $D(y,\hat{y})$ measures the minimum number of single-character insertions, deletions, or substitutions required to convert $\hat{y}$ into $y$, thereby capturing both semantic and formatting discrepancies. This reward is bounded within $[0,1]$, with higher values indicating better alignment between prediction and reference. Meanwhile, the reference output $y$ is synthesized through two data generation pipelines proposed in this work. It is produced with rigorous rule-based filtering and consistency validation, and serves as a high-quality surrogate for ground-truth annotations in evaluating the quality of the model output.
2. Count Reward ($R_{\mathrm{count}}$)
To encourage accurate paragraph segmentation, let $N_{Y}$ and $N_{\hat{Y}}$ be the numbers of reference and predicted paragraphs. We define:
$R_{\text{count}}=1-\frac{|N_{Y}-N_{\hat{Y}}|}{N_{Y}},$
which penalizes missing or spurious paragraphs.
3. Order Reward ($R_{\mathrm{order}}$)
We measure sequence-level fidelity by counting pairwise inversions $D_{\mathrm{order}}$ between reference and predicted paragraphs. with $\max_{\mathrm{inv}}=N_{Y}(N_{Y}-1)/2$, we set:
$R_{\text{order}}=1-\frac{D_{\text{order}}}{\max_{\text{inv}}},$
rewarding preservation of the original reading order.
The final multi-aspect reward is a weighted combination of these three components. Specifically, we begin by applying the Hungarian algorithm [20] to establish the optimal one-to-one matching between predicted and ground-truth segments, identifying both pairings and their relative order. Based on the matched segment count, we compute the count reward to reflect alignment in the number of segments. Using the relative sequence of matched pairs, we calculate the order reward to measure structural consistency. On top of these matchings, we compute the edit reward by averaging the edit similarities of each matched segment pair. Combining these terms yields the final reward:
$R_{\text{Multi-Aspect}}=R_{\text{dist}}+R_{\text{count}}+R_{\text{order}}$
This multi-aspect design balances content fidelity with structural correctness and order preservation, providing rich supervision for end-to-end document parsing.
3 Experiments
We fine-tune the Qwen2.5-VL-7B model using GRPO within a distributed training setup based on Verl [45, 66], utilizing 8 A100 GPUs (80GB). Throughout our experiments, we set the KL coefficient $\beta=1.0\times 10^{-2}$. And for each problem instance, we sample 8 responses, each with a maximum length of 8192 tokens and a temperature of 1.0. Both the rollout batch size and the global batch size are set to 128. The actor model is updated using the AdamW optimizer with parameters $(\beta_{1}=0.9,\beta_{2}=0.99)$ and a learning rate $1.0\times 10^{-6}$. The model is trained for 1.0 epoch for all experiments.
[…]
EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH
Based on Pipeline Tools
MinerU [52] 0.061 0.215 0.278 0.577 57.3 42.9 78.6 62.1 0.18 0.344 0.079 0.292 0.15 0.357
Marker  [37] 0.080 0.315 0.530 0.883 17.6 11.7 67.6 49.2 0.619 0.685 0.114 0.340 0.336 0.556
Mathpix 0.105 0.384 0.306 0.454 62.7 62.1 77.0 67.1 0.243 0.320 0.108 0.304 0.191 0.365
Docling [30] 0.416 0.987 0.999 1.000 61.3 25.0 0.627 0.810 0.313 0.837 0.589 0.909
Pix2Text [12] 0.138 0.356 0.276 0.611 78.4 39.6 73.6 66.2 0.584 0.645 0.281 0.499 0.320 0.528
Unstructured-0.17.2 0.198 0.481 0.999 1.000 - - - - 1.000 0.998 0.145 0.387 0.586 0.716
OpenParse-0.7.0 0.681 0.974 0.996 1.000 - - 64.8 27.5 0.284 0.639 0.595 0.641 0.646 0.814
Based on Expert VLMs
GOT-OCR [59] 0.189 0.315 0.360 0.528 74.3 45.3 53.2 47.2 0.459 0.520 0.141 0.280 0.287 0.411
Nougat [6] 0.365 0.998 0.488 0.941 15.1 16.8 39.9 0.0 0.572 1.000 0.382 0.954 0.452 0.973
Mistral OCR 0.072 0.325 0.318 0.495 64.6 45.9 75.8 63.6 0.600 0.650 0.083 0.284 0.268 0.439
OLMOCR-sglang 0.097 0.293 0.455 0.655 74.3 43.2 68.1 61.3 0.608 0.652 0.145 0.277 0.326 0.469
SmolDocling-256M 0.262 0.838 0.753 0.997 32.1 55.1 44.9 16.5 0.729 0.907 0.227 0.522 0.493 0.816
Based on General VLMs
GPT-4o [1] 0.144 0.409 0.425 0.606 72.8 42.8 72.0 62.9 0.234 0.329 0.128 0.251 0.233 0.399
Qwen2-VL-72B [55] 0.096 0.218 0.404 0.487 82.2 61.2 76.8 76.4 0.387 0.408 0.119 0.193 0.252 0.327
InternVL2-76B [8] 0.353 0.290 0.543 0.701 67.4 44.1 63.0 60.2 0.547 0.555 0.317 0.228 0.440 0.443
Qwen2.5-VL-7B [4] 0.222 0.236 0.394 0.587 68.4 45.1 71.5 75.7 0.465 0.294 0.246 0.207 0.332 0.331
InternVL3-8B [75] 0.315 0.345 0.714 0.729 59.4 72.3 59.0 71.5 0.352 0.211 0.324 0.257 0.426 0.385
Based on Reinforcement Learning
[…]
Table 2: Comprehensive evaluation of document parsing algorithms on OmniDocBench: performance metrics for text, formula, table, and reading order extraction, with overall scores derived from ground truth comparisons.
3.1 Evaluation and Metrics
OmniDocBench [36]
We introduces an end-to-end evaluation pipeline to ensure fair and unified assessment of document parsing models. It processes model-generated markdown by cleaning, extracting key elements (tables, formulas, code), standardizing inline formulas, and recording positions for reading order evaluation. This comprehensive pipeline enables reliable, multi-level evaluation across diverse document types. We evaluate different content types using tailored metrics :
  • Pure Text: Normalized edit distance (NED) is computed and averaged at the sample level to obtain the final scores.
  • Tables: Tables are converted to HTML, then evaluated using TEDS [71] and NED.
  • Formulas: Evaluation is based on CDM [51], NED, and BLEU.
  • Reading Order: Evaluated by NED, considering only text and excluding other components.
Fox [26]
A multilingual benchmark for fine-grained, text-centric document understanding, covering 9 sub-tasks such as OCR, translation, summarization, layout analysis, and captioning. It includes 212 bilingual pages (112 English, 100 Chinese) with diverse layouts, evaluated using task-specific metrics like NED, F1, BLEU, METEOR, and ROUGE-L.
PubTabNet [71]
A widely used benchmark for table recognition, containing 500,777 training and 9,115 validation images with diverse scientific table structures. Evaluation is conducted on the validation set.
FinTabNet [70]
Focused on financial documents, this dataset includes 112,000 single-page scanned documents, with 92,000 cropped training images and 10,656 for testing. It features dense layouts and detailed annotations for both structure and content evaluation.
olmOCR-Bench [39]
This is a benchmark developed to automatically and reliably evaluate document-level OCR performance across a wide range of tools. Unlike traditional evaluation metrics such as edit distance—which may penalize valid variations or fail to capture critical semantic errors—olmOCR-Bench focuses on verifying simple, unambiguous, and machine-checkable “facts” about each document page, similar to unit tests. For instance, it checks whether a specific sentence appears exactly in the OCR output. The benchmark operates directly on single-page PDFs to preserve digital metadata, which can be beneficial for certain OCR systems, and to maintain the integrity of the original document format. Designed for flexibility, olmOCR-Bench supports outputs in Markdown or plain text, allowing for seamless evaluation of both open-source and custom OCR pipelines.
Model Language Table Frame Type Special Situation (+/-) Overall $\uparrow$
EN ZH Mixed Full Omission Three Zero MergeCell Form. Colorful Rotate
PaddleOCR [22] 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
RapidTable [40] 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
StructEqTable [74] 72.8 75.9 83.4 72.9 76.2 76.9 88.0 64.5/81.0 69.2/76.6 72.8/76.4 30.5/76.2 75.8
GOT-OCR [59] 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
Qwen2-VL-7B [55] 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
InternVL2-8B [8] 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
Qwen2.5-VL-7B [55] 87.4 80.7 93.5 86.4 85.1 84.1 88.7 77.5/89.8 82.1/87.2 77.1/87.5 56.5/86.0 85.5
InternVL3-8B [75] 79.5 86.0 91.7 85.5 80.7 83.9 85.9 71.9/90.9 74.0/86.7 82.1/85.3 12.6/85.5 84.3
[…]
Table 3: Component-level Table Recognition evaluation on OmniDocBench table subset. (+/-) means with/without special situation.
3.2 Main Results
In this section, we evaluate the model’s performance in document parsing, table recognition, and Document-level OCR. In this work, we ensure that the test data for each benchmark undergoes rigorous text similarity filtering to prevent any overlap with the training data.
Overall Evaluation on OmniDocBench As shown in Table 2, pipeline-based methods such as MinerU [52] and Mathpix achieve superior performance across individual sub-tasks including text recognition, formula recognition, and table recognition. Meanwhile, general-purpose vision-language models like Qwen2-VL and GPT-4o also demonstrate competitive results. Notably, most methods perform better on English pages compared to Chinese pages, reflecting language-dependent challenges. In contrast, our proposed Infinity-Parser-7B achieves a more balanced performance across all sub-tasks and languages, setting new SOTA results in table recognition and maintaining competitive accuracy in text, formula, and reading order extraction. This highlights the advantage of reinforcement learning with multi-aspect rewards in enabling robust, end-to-end document parsing.
Table Recognition Evaluation To evaluate the model’s generalization ability, we introduce task-specific test cases. In Table 4, we compare Infinity-Parser-7B with end-to-end table recognition models on PubTabNet and FinTabNet using the TEDS metric, which evaluates both structure and content. We also report TEDS-S for structure-only assessment. The evaluation results for InternVL3, Qwen2.5-VL, and GPT-4o were generated through our standardized benchmarking pipeline. Infinity-Parser-7B achieves the highest TEDS-S and TEDS scores on both datasets.
Model PubTabNet [71] FinTabNet [70]
TEDS-S TEDS TEDS-S TEDS
EDD [72] 89.9 88.3 90.6 -
OmniParser [49] 90.45 88.83 91.55 89.75
InternVL3-8B [75] 87.48 83.02 86.73 84.01
InternVL3-78B [75] 89.63 82.11 92.51 89.21
Qwen2.5-VL-7B [3] 86.78 81.60 87.46 82.58
Qwen2.5-VL-72B [3] 87.91 84.39 87.13 82.90
GPT-4o [1] 86.16 76.53 87.00 83.96
[…]
Table 4: Comparisons of end-to-end table recognition methods on PubTabNet and FinTabNet.
Table 3 summarizes the performance of various models on the OmniDocBench table subset, evaluated along three dimensions: language diversity, table frame types, and special layout conditions. Notably, Infinity-Parser-7B achieves the best overall performance with an impressive score of 86.4, outperforming all other models across most individual metrics. It leads in nearly every category, including mixed-language settings (94.8), complex frame layouts (e.g., omission and three-line formats), and challenging special situations such as merged cells, formulas, and rotations. This demonstrates its strong generalization ability and robustness across diverse and noisy table formats.
Document-level OCR Evaluation As shown in Table 5, Infinity-Parser-7B achieves the best edit distance among all evaluated models, scoring 0.023 for English and 0.017 for Chinese, outperforming all baselines including large-scale models such as Qwen-VL-Max [2] and Fox [26]. This result suggests that Infinity-Parser-7B can reconstruct dense document text with exceptional precision and minimal character-level errors.
[…]
en zh en zh en zh en zh en zh en zh
UReader [67] 7B 0.344 - 0.296 - 0.469 - 0.103 - 0.287 - 0.718 -
LLaVA-NeXT [27] 34B 0.647 - 0.573 - 0.881 - 0.478 - 0.582 - 0.430 -
InternVL-ChatV1.5 [7] 26B 0.751 0.816 0.698 0.784 0.917 0.866 0.568 0.622 0.663 0.717 0.393 0.265
Nougat [5] 250M 0.745 - 0.720 - 0.809 - 0.665 - 0.761 - 0.255 -
TextMonkey [28] 7B 0.821 - 0.778 - 0.906 - 0.671 - 0.762 - 0.265 -
DocOwl1.5 [14] 7B 0.862 - 0.835 - 0.962 - 0.788 - 0.858 - 0.258 -
Vary [57] 7B 0.918 0.952 0.906 0.961 0.956 0.944 0.885 0.754 0.926 0.873 0.092 0.113
Vary-toy [58] 1.8B 0.924 0.914 0.919 0.928 0.938 0.907 0.889 0.718 0.929 0.832 0.082 0.142
Qwen-VL-Plus [2] - 0.931 0.895 0.921 0.903 0.950 0.890 0.893 0.684 0.936 0.828 0.096 0.121
[…]
Fox [26] 1.8B 0.952 0.954 0.957 0.964 0.948 0.946 0.930 0.842 0.954 0.908 0.046 0.061
GOT [60] 580M 0.972 0.980 0.971 0.982 0.973 0.978 0.947 0.878 0.958 0.939 0.035 0.038
[…]
Table 5: Performance comparison of dense English and Chinese OCR on document-level pages.
Models Book Slides Financial Report Textbook Exam Paper Magazine Academic Papers Notes Newspaper Overall $\downarrow$
Based on Pipeline Tools
MinerU 0.055 0.124 0.033 0.102 0.159 0.072 0.025 0.984 0.171 0.206
Marker 0.074 0.34 0.089 0.319 0.452 0.153 0.059 0.651 0.192 0.274
Mathpix 0.131 0.22 0.202 0.216 0.278 0.147 0.091 0.634 0.69 0.3
Based on Expert VLMs
GOT-OCR 0.111 0.222 0.067 0.132 0.204 0.198 0.179 0.388 0.771 0.267
Nougat 0.734 0.958 1.000 0.820 0.930 0.83 0.214 0.991 0.871 0.806
Based on General VLMs
GPT-4o 0.157 0.163 0.348 0.187 0.281 0.173 0.146 0.607 0.751 0.316
Qwen2-VL-72B 0.096 0.061 0.047 0.149 0.195 0.071 0.085 0.168 0.676 0.179
InternVL2-76B 0.216 0.098 0.162 0.184 0.247 0.150 0.419 0.226 0.903 0.3
Qwen2.5-VL-7B 0.222 0.131 0.194 0.268 0.203 0.230 0.195 0.249 0.394 0.230
InternVL3-8B 0.311 0.233 0.320 0.222 0.238 0.157 0.438 0.268 0.726 0.328
Based on Reinforcement Learning
[…]
Table 6: End-to-end text recognition performance on OmniDocBench: evaluation using edit distance across 9 PDF page types. We compare with Mathpix, MinerU [52], Marker [37], GOT-OCR [59], Nougat [6], GPT-4o [1], Qwen2-VL-72B [55], InternVL2-76B [8], Qwen2.5-VL-7B [4], InternVL3-8B [75].
Diverse Page Types Evaluation To further investigate model behavior across diverse document types, we evaluated text recognition performance on nine distinct page categories. As shown in Table 6, pipeline-based systems such as MinerU [52] and Mathpix achieved strong results on structured formats like academic papers and financial reports. General-purpose vision-language models (VLMs) demonstrated better generalization on less formal page types, including presentation slides and handwritten notes. However, for challenging formats such as newspapers, most VLMs underperformed, while pipeline tools maintained relatively lower error rates. Notably, our proposed Infinity-Parser-7B achieved consistently low edit distances across all document types, outperforming both pipeline-based systems and general-purpose VLMs in overall accuracy. This highlights the robustness and adaptability of our reinforcement learning approach across diverse and complex document layouts.
Model Anchor ArXiv Math Tables Scans H&F MultiCol TinyText Base Overall
GOT OCR 52.7 52.0 0.2 22.1 93.6 42.0 29.9 94.0 48.3
Marker v1.6.2 24.3 22.1 69.8 24.3 87.1 71.0 76.9 99.5 59.4
MinerU v1.3.10 75.4 47.4 60.9 17.3 96.6 59.0 39.1 96.6 61.5
Mistral OCR API 77.2 67.5 60.6 29.3 93.6 71.3 77.1 99.4 72.0
GPT-4o 51.5 75.5 69.1 40.9 94.2 68.9 54.1 96.7 68.9
GPT-4o 53.5 74.5 70.0 40.7 93.8 69.3 60.6 96.8 69.9
Gemini Flash 2 32.1 56.3 61.4 27.8 48.0 58.7 84.4 94.0 57.8
Gemini Flash 2 54.5 56.1 72.1 34.2 64.7 61.5 71.5 95.6 63.8
Qwen 2 VL 19.7 31.7 24.2 17.1 88.9 8.3 6.8 55.5 31.5
Qwen 2.5 VL 63.1 65.7 67.3 38.6 73.6 68.3 49.1 98.3 65.5
olmOCR v0.1.68 72.1 74.7 71.5 43.7 91.6 78.5 80.5 98.1 76.3
olmOCR v0.1.68 75.6 75.1 70.2 44.5 93.4 79.4 81.7 99.0 77.4
[…]
Table 7: Performance comparison on the olmOCR [39] benchmark across diverse document structures. ArXiv: scientific articles. Math: scanned mathematical content. Scans: historical scanned documents. H&F: headers and footers. MultiCol: multi-column layout. TinyText: long lines or extremely small text. Base: base text recognition accuracy. Higher is better. The Anchor column indicates whether external PDF metadata (e.g., layout hints) is used to assist layout prediction.
Document-Level OCR Evaluation on olmOCR-Bench Table 7 reports performance on the olmOCR-Bench benchmark, which evaluates document-level OCR across diverse layouts and domains. Infinity-Parser-7B achieves the highest overall score (79.1), followed closely by olmOCR v0.1.68 (Anchored) (77.4), both demonstrating strong performance in complex categories like multi-column layouts and scanned math content. The results highlight the effectiveness of anchored prompting, with anchored versions of models (e.g., GPT-4o, olmOCR) significantly outperforming their non-anchored counterparts—especially on tables and old scans. This underscores the importance of layout-aware extraction techniques. In contrast, traditional pipelines like Marker and GOT OCR lag behind in structural accuracy, reinforcing the value of modern VLM-based approaches in high-fidelity PDF understanding.
3.3 Ablation Study
We perform ablation experiments to evaluate the individual contributions of our three core design choices: (1) data quality verification and (2) multi-aspect rewards. We report all ablation results using two primary evaluation metrics: OverallEdit and OverallCat.. OverallEdit represents the average edit-based overall score across English and Chinese pages, as shown in Table 2. In contrast, OverallCat. reflects the mean category-level performance across nine types of scanned documents pages, following the same evaluation setting as Table 10 in the Appendix. To save computational resources, the ablation studies are conducted on a randomly selected subset of 22K samples, rather than using the full 55K dataset as in the final results.
Effect of Data Validity Table 8 highlights two important findings regarding data effectiveness and the limitations of the SFT paradigm. First, increasing the size of training data improves performance in terms of OverallCat., which drops significantly from 0.257 at 5k to 0.166 at 22k, confirming the value of larger datasets in enhancing content categorization accuracy. However, the OverallEdit metric shows a non-monotonic trend: while initial gains are observed from 5k to 10k, performance begins to degrade as the dataset size increases further. This suggests that beyond a certain point, the SFT paradigm struggles to maintain structural precision, possibly due to the lack of layout-aware supervision, revealing its limitations in scaling for layout-sensitive document parsing.
As shown in Table 9, we evaluate the impact of data scale on RL training with layout-aware rewards. The zero-shot baseline performs worst, highlighting the importance of training data. From 5k to 15k, category error steadily decreases and edit distance improves, indicating better generalization. At 22k, performance peaks, though minor edit distance fluctuations suggest possible noise. These results show that scaling data enhances RL’s structural modeling capabilities.
Method Data Size OverallEdit $\downarrow$ OverallCat. $\downarrow$
Zero Shot 0.346 0.259
SFT 5k 0.332 0.257
SFT 10k 0.353 0.167
SFT 15k 0.374 0.179
SFT 22k 0.375 0.166
Table 8: Results with different data sizes.
Method Data Size OverallEdit $\downarrow$ OverallCat. $\downarrow$
Zero-Shot - 0.346 0.259
RL 5k 0.275 0.168
RL 10k 0.310 0.153
RL 15k 0.324 0.147
RL 22k 0.260 0.123
Table 9: Results with different data types.
Effect of Data Construction Methods Table 10 demonstrates that synthetic data serves as an effective complement to real-world data for document parsing. Under both SFT and RL settings, synthetic data consistently achieves lower scores in OverallEdit and OverallCat., indicating better structural alignment and parsing accuracy. Notably, combining real and synthetic data under the RL setup yields the best overall performance,
Method Data Type OverallEdit $\downarrow$ OverallCat. $\downarrow$
SFT Real 0.337 0.279
SFT Synth 0.290 0.214
RL Real 0.285 0.164
RL Synth 0.262 0.156
RL Synth+Real 0.273 0.155
Table 10: Results with different data sizes.
especially in OverallCat. (0.155), suggesting that the structural precision of synthetic documents enhances the diversity and complexity offered by real-world samples. These results highlight the importance of leveraging both data sources to achieve robust and layout-aware document understanding.
Method Edit Dist. Reward Count Reward Order Reward SFT RL OverallEdit $\downarrow$ OverallCat. $\downarrow$
Zero Shot - - - - - 0.346 0.259
SFT - - - 22k - 0.375 0.166
SFT + RL - - 5k 17k 0.345 0.244
Zero + RL - - - 22k 0.287 0.156
Zero + RL - - 22k 0.280 0.141
Zero + RL - 22k 0.260 0.123
Table 11: Results under different reward designs.
Effect of Multi-Aspect Rewards Table 11 shows that reinforcement learning consistently outperforms supervised fine-tuning, particularly under layout-aware reward designs. Compared to the SFT baseline (0.375 / 0.166), the RL method with distance-based reward (Zero + $R_{\text{dist}}$) achieves better OverallEdit (0.287) and comparable OverallCat. (0.156). Further introducing count and order rewards leads to significant improvements in structural consistency, with Zero + $R_{\text{dist}}+R_{\text{count}}$ achieving 0.280 / 0.141, and the full reward setting reaching 0.260 / 0.123. These results highlight RL’s ability to incorporate structural inductive biases via reward design, enabling better alignment with task-specific goals and reducing overfitting to token-level patterns. The improvements across reward configurations confirm the effectiveness of multi-aspect rewards in enhancing layout-aware generalization.
(a) Comparison of four Markdown generation results on a single case, illustrating progressive improvement from direct inference to full reward integration.
(b) Markdown extraction from a PowerPoint-style PDF slide. The figure includes the original slide (a), human-annotated ground truth (b), and outputs from MinerU, GPT-4o, and Infinity-Parser (c–e). MinerU and GPT-4o introduce spacing and formatting errors. In contrast, Infinity-Parser accurately preserves dialogue structure and layout fidelity.
(a) Comparison of Markdown extraction from academic literature. The figure shows the original PDF (a), ground truth annotations (b), and extraction results from MinerU, GPT-4o, and Infinity-Parser (c–e). Infinity-Parser produces the most accurate output.
(b) Comparison of Markdown extraction from a book-style PDF. GPT-4o introduces multiple errors; MinerU retains unnecessary line breaks; Infinity-Parser produces clean and accurate output.
(a) Markdown extraction from a colorful textbook-style PDF. The figure displays the original illustrated page (a), ground truth annotations (b), and outputs from MinerU, GPT-4o, and Infinity-Parser (c–e). MinerU and GPT-4o introduce issues such as redundant content, missed spacing, and incorrect formatting. In contrast, Infinity-Parser accurately identifies section titles and body text, preserving the structure and semantics of the original content.
(b) Markdown extraction from an exam-style PDF. The figure presents the original exam document (a), ground truth annotations (b), and the extraction results from MinerU, GPT-4o, and Infinity-Parser (c–e). GPT-4o and MinerU both introduce redundant text and formatting errors, such as incorrect title levels and misplaced content. In contrast, Infinity-Parser accurately captures the heading hierarchy and structured list format.
(a) Markdown extraction from a magazine-style PDF. The figure shows the original visually-rich page (a), ground truth annotations (b), and results from MinerU, GPT-4o, and Infinity-Parser (c–e). GPT-4o suffers from significant formatting errors, while Infinity-Parser accurately preserves the structural hierarchy and formatting.
(b) Markdown extraction from a newspaper-style PDF. This figure presents the original densely formatted page (a), ground truth annotations (b), and the results from MinerU, GPT-4o, and Infinity-Parser (c–e). Both MinerU and GPT-4o exhibit issues such as incorrect title levels, redundant content, and format errors. In contrast, Infinity-Parser accurately identifies the main title, maintains structural formatting, and preserves content integrity.
3.4 Result Analysis
Figure 5(a) illustrates a progressive improvement in Markdown generation quality across different training strategies. The zero-shot model fails to capture key structural elements, omitting titles and producing redundant or incomplete content. With SFT, the model better identifies section headers and general layout but still suffers from symbol-level errors and repeated outputs. In contrast, the layout-aware RL model demonstrates the most accurate and coherent result, successfully preserving the document hierarchy and eliminating redundancy. This highlights the effectiveness of layout-aware rewards in guiding the model toward semantically and structurally faithful document parsing. Infinity-Parser consistently outperforms baselines across a wide range of document types, including academic papers, books, textbooks, exams, magazines, government notices, newspapers, and slides. It excels in structure parsing, content recognition, and formatting accuracy, demonstrating strong generalization and layout robustness (Figures 5(b) - 8(b)).
4 Related Work
4.1 Reinforcement Learning for Language Models
Recent advancements in Large Language Models (LLMs) such as OpenAI’s GPT series [35]), DeepSeek-R1 [11], and Gemini [47] have highlighted the significant potential of Reinforcement Learning (RL) in enhancing their reasoning capabilities. This RL paradigm has been successfully extended to other domains demanding sophisticated reasoning, including code generation [23, 68], autonomous tool utilization [41, 53], and information retrieval [33]. Similarly, RL has demonstrated its efficacy in the domain of Visual Language Models (VLMs), including precise object counting [38], nuanced visual perception [29], and complex multimodal reasoning (e.g., Pixel Reasoner [46], VL-Rethinker [54], Vision-R1 [16]). These pioneering works have predominantly relied on binary outcome rewards to guide RL training. Complementary to these efforts, our work demonstrates the effectiveness of incorporating layout-aware and layout-based rewards for document parsing, offering a more granular and contextually relevant feedback mechanism.
4.2 VLM-based Document Parsing
Recent advancements in document understanding and optical character recognition (OCR) have highlighted their importance as critical benchmarks for evaluating the perceptual capabilities of vision-language models (VLMs). By incorporating large-scale OCR corpora during pretraining, models such as GPT-4o [1] and Qwen2-VL [2] have achieved competitive performance on document content extraction tasks. Building upon these foundations, the emergence of VLMs has further accelerated the progress of end-to-end document parsing, giving rise to a range of models such as Donut [6], Nougat [5], Kosmos-2.5 [31], Vary [61], mPLUG-DocOwl [15], Fox [25], and GOT [59]. These models have continued to improve their understanding of visual layouts and textual content by leveraging advancements in visual encoders [10], language decoders [2], and data construction pipelines. Despite the success of these VLM-based approaches in enabling end-to-end document parsing, they still face generalization challenges on downstream layout parsing tasks [52, 59]. To address this issue, we propose leveraging reinforcement learning to provide a more effective training paradigm that better aligns with the demands of document parsing.
5 Conclusion
In this work, we presented layoutRL, the first end-to-end reinforcement-learning framework that makes document parsers explicitly layout-aware by optimizing a multi-aspect reward—combining normalized edit-distance, paragraph-count accuracy, and reading-order preservation. To support this training paradigm, we released Infinity-Doc-55K, a 55K document corpus blending high-fidelity synthetic scanned document parsing data with expert-filtered real-world samples, providing the large-scale, precise supervision end-to-end models require. We implement the proposed method as a vision-language-model–based parser, Infinity-Parser. Powered by layoutRL, it achieves new state-of-the-art performance among end-to-end models on English and Chinese tasks, including OCR, table and formula extraction, and reading order detection. By releasing our code, dataset, and trained model, we hope to catalyze further advances in robust, reliable document understanding.